OpenStudy (anonymous):

Anybody can help? Find the center and the foci of the ellipse given by the equation -y^2+10y-4=9x^2+12x. At what points, if any, does this ellipse intersect the line y= 2(x+10)?

OpenStudy (anonymous):

\[(x-h)^{2}/a ^{2} + (y-k)^{2}/b ^{2} =1\] try to get equation in these form

OpenStudy (anonymous):

I know this equation. But I cant get the answer at the end :(

OpenStudy (dumbcow):

use substitution replace y with 2(x+10) in ellipse equation

OpenStudy (anonymous):

see.... @AliceMath99 This equation may be rewritten as: 9x2 + Y2+ 12X -10Y +4=0 to find the center we have h= 12/2*9 = (2/3) = Coefficient of x divided by 2 times coefficient X2 k= 10/2=5 coefficient of Y divided by two times coefficient of Y2 The center is at ((2/3),5) the foci are at: a=√1=1 b= √9 =3 then the distance between the foci is 2c=2*√(b2-a2) = 4√2 Then the foci are located at: ((2/3), (5+4√2)) and ((2/3), (5-4√2)) Please do the math...... Hope it helps!!